# Enabling Gigantic MIMO Beamforming

### v1

@@ Abstract
In our previous work, we have introduced a microwave linear analog computer (MiLAC) as an analog computer that processes microwave signals linearly, demonstrating its potential to reduce the computational complexity of specific signal processing tasks. In this paper, we extend these benefits to wireless communications, showcasing how MiLAC enables gigantic multiple-input multiple-output (MIMO) beamforming entirely in the analog domain. MiLAC-aided beamforming can implement regularized zero-forcing beamforming (R-ZFBF) at the transmitter and minimum mean square error (MMSE) detection at the receiver, while significantly reducing hardware costs by minimizing the number of radio-frequency (RF) chains and only relying on low-resolution analog-to-digital converters (ADCs) and digital-to-analog converters (DACs). In addition, it eliminates per-symbol operations by completely avoiding digital-domain processing and remarkably reduces the computational complexity of R-ZFBF, which scales quadratically with the number of antennas instead of cubically. Numerical results show that it can perform R-ZFBF with a computational complexity reduction of up to 7400 times compared to digital beamforming.
@@
Analog computing, gigantic multiple-input multiple-output (MIMO), minimum mean square error (MMSE), zero-forcing (ZF).
@@ I Introduction
Future wireless systems are expected to require a significantly larger number of antennas to support higher data rates and simultaneously serve a larger number of users. While massive multiple-input multiple-output (MIMO) systems typically deploy 64 or more antennas [1, 2], this number could reach several thousands in ultra-massive or gigantic MIMO systems, which have been recently proposed to meet the stringent requirements of 6G networks [3]. With such a large number of antennas, digital beamforming faces two critical challenges in terms of hardware and computational complexity. First, digital beamforming typically requires a dedicated radio-frequency (RF) chain per antenna element, each including analog-to-digital converters (ADCs)/digital-to-analog converters (DACs) and mixers, which are costly and power-hungry components. Second, as the number of antennas increases, the data volumes to be processed in real-time grow accordingly. For example, to precode the transmitted symbol vector at the transmitter (or combine the received symbol vector at the receiver) a matrix-vector product is required on a per-symbol basis. These two challenges result in high hardware costs and computational complexity when digital beamforming is applied in massive or gigantic MIMO systems. For these reasons, several alternatives to digital beamforming have been explored to enable beamforming operations fully or partially in the analog domain.
@@
To address the high hardware cost associated with digital beamforming, alternative beamforming strategies have been proposed. Analog beamforming steers the beam pattern by solely requiring a single RF chain [4]. This is achieved by connecting the RF chain to the antennas via phase shifters or time delay elements, which impose a constant modulus constraint on the beamforming weights. To balance hardware cost and flexibility by bridging analog and digital beamforming, hybrid analog-digital beamforming techniques have been widely investigated [4, 5, 6]. Another method to efficiently perform beamforming in the analog domain is given by reconfigurable intelligent surface (RIS) [7, 8]. This technology has recently emerged enabling the control of the wireless channel through surfaces made of multiple passive elements with reconfigurable scattering properties. Beamforming in the analog domain can be achieved through RIS by deploying a RIS in close proximity to an active transceiving device equipped with a reduced number of antennas. In this way, the effective radiation pattern of the active device can be optimized by reconfiguring the RIS elements [9, 10, 11]. Two extensions of this technology have been proposed to achieve additional flexibility over conventional RIS, such as beyond diagonal RIS (BD-RIS) [12] and stacked intelligent metasurface (SIM) [13]. The additional flexibility provided by BD-RIS and the multiple layers in SIM has shown increased performance over just deploying a single conventional RIS.
@@
Numerous technologies have been proposed to reduce the hardware cost of future wireless transceivers by executing beamforming in the analog domain [4]-[13]. However, the use of analog beamforming strategies to alleviate the computational complexity of future wireless networks has received limited attention [14]. In this paper, we address this gap by proposing a novel beamforming strategy enabled by the concept of microwave linear analog computer (MiLAC), recently introduced in [15]. This beamforming strategy, which operates entirely in the analog domain, reduces the hardware cost by minimizing the number of required RF chains and relying only on low-resolution ADCs/DACs. By leveraging the computational capabilities of MiLACs, it also minimizes the computational complexity required for beamforming.
@@
The contributions of this paper can be summarized as follows. First, we demonstrate that a MiLAC can efficiently compute regularized least squares (RLS), with significantly reduced computational complexity compared to conventional digital computing. Second, we propose MiLAC-aided beamforming as a new beamforming strategy enabled by MiLAC scalable to gigantic MIMO systems. MiLAC-aided beamforming allows us to perform regularized zero-forcing beamforming (R-ZFBF) at the transmitter and minimum mean square error (MMSE) detection at the receiver in the analog domain, with important gains in terms of both hardware and computational complexity. Third, we provide numerical results to assess the performance of MiLAC-aided beamforming in terms of sum rate, confirming that it can achieve the same performance as digital beamforming. We also assess the benefits of MiLAC-aided beamforming over digital beamforming in terms of computational complexity, showing that it can perform R-ZFBF with a gain in computational complexity of $7.4\times 10^{3}$ times. Thus, in a gigantic MIMO system with thousands of antennas, MiLAC-aided beamforming achieves the same performance as digital R-ZFBF at just $1/7400$th of the computational cost. From a different perspective, a 4096-antenna MiLAC serving 4096 users via R-ZFBF requires the same computational complexity as a 256-antenna digital transmitter serving 256 users. This makes MiLAC-aided beamforming a game changer for gigantic MIMO, opening the door to radically new transceiver architectures that can massively scale the number of antennas and users.
Notation: Vectors and matrices are denoted with bold lower and bold upper letters, respectively. Scalars are represented with letters not in bold font. $\mathbf{a}^{T}$, $\mathbf{a}^{H}$, $[\mathbf{a}]_{i}$, and $\|\mathbf{a}\|_{2}$ refer to the transpose, conjugate transpose, $i$th element, and $l_{2}$-norm of a vector $\mathbf{a}$, respectively. $\mathbf{A}^{T}$, $\mathbf{A}^{H}$, $[\mathbf{A}]_{i,k}$, $[\mathbf{A}]_{:,k}$, and $\|\mathbf{A}\|_{F}$ refer to the transpose, conjugate transpose, $(i,k)$th element, $k$th column, and Frobenius norm of a matrix $\mathbf{A}$, respectively. $\mathbb{R}$ and $\mathbb{C}$ denote the real and complex number sets, respectively. $j=\sqrt{-1}$ denotes the imaginary unit. $\mathbf{I}_{N}$ and $\mathbf{0}_{N}$ denote the identity matrix and the all-zero matrix with dimensions $N\times N$, respectively, and $\mathbf{0}_{M\times N}$ denote the all-zero matrix with dimensions $M\times N$.
@@ II Microwave Linear Analog Computer
In this section, we briefly review the concept of MiLAC, defined as an analog computer that linearly processes microwave signals, as introduced in [15]. In general, a MiLAC can be seen as a $P$-port reconfigurable microwave network, which can be characterized by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$. We assume this microwave network to be made of $P^{2}$ tunable admittance components interconnecting each port to ground and to all the other ports, such that the microwave network can be reconfigured to assume any arbitrary admittance matrix $\mathbf{Y}$. The admittance component interconnecting port $k$ to ground is denoted as $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, while the admittance component interconnecting port $k$ to port $i$ is denoted as $Y_{i,k}\in\mathbb{C}$, for $\forall i\neq k$. The input signal $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$ is applied by $N$ voltage generators with series admittance $Y_{0}$ to the first $N$ ports of the MiLAC, with $N<P$, as represented in Fig. 1. The output signal is read on all $P$ ports, where $P=N+M$ with $M$ denoting the number of ports with no input, and is denoted by $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$, as shown in Fig. 1. For convenience of representation, we introduce the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as $\mathbf{v}_{1}=[v_{1},\ldots,v_{N}]^{T}\in\mathbb{C}^{N\times 1}$ and $\mathbf{v}_{2}=[v_{N+1},\ldots,v_{p}]^{T}\in\mathbb{C}^{M\times 1}$, such that $\mathbf{v}=[\mathbf{v}_{1}^{T},\mathbf{v}_{2}^{T}]^{T}$.
It has been shown in [15] that the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ can be expressed as a function of the admittance matrix $\mathbf{Y}$ and the input $\mathbf{u}$ by introducing a matrix $\mathbf{P}\in\mathbb{C}^{P\times P}$ as
@@
for $i,k=1,\ldots,P$, as discussed in [15]. Remarkably, a MiLAC can efficiently compute (3)-(4) and (5)-(6) in the analog domain, while they would require computationally expensive matrix-matrix products and matrix inversion operations if digitally computed.
TABLE I: MiLAC output $\mathbf{v}_{2}$ for different $\mathbf{P}$ with $\mathbf{P}_{11}$ invertible.

### v2

# Analog Computing for Signal Processing and Communications – Part II: Toward Gigantic MIMO Beamforming

@@ Abstract
Analog-domain operations offer a promising solution to accelerating signal processing and enabling future multiple-input multiple-output (MIMO) communications with thousands of antennas. In Part I of this paper, we have introduced a microwave linear analog computer (MiLAC) as an analog computer that processes microwave signals linearly, demonstrating its potential to reduce the computational complexity of specific signal processing tasks. In Part II of this paper, we extend these benefits to wireless communications, showcasing how MiLAC enables gigantic MIMO beamforming entirely in the analog domain. MiLAC-aided beamforming enables the maximum flexibility and performance of digital beamforming, while significantly reducing hardware costs by minimizing the number of radio-frequency (RF) chains and only relying on low-resolution analog-to-digital converters (ADCs) and digital-to-analog converters (DACs). In addition, it eliminates per-symbol operations by completely avoiding digital-domain processing and remarkably reduces the computational complexity of zero-forcing (ZF), which scales quadratically with the number of antennas instead of cubically. It also processes signals with fixed matrices, e.g., the discrete Fourier transform (DFT), directly in the analog domain. Numerical results show that it can perform ZF and DFT with a computational complexity reduction of up to $1.5\times 10^{4}$ and $4.0\times 10^{7}$ times, respectively, compared to digital beamforming.
@@
Analog beamforming, analog computing, massive multiple-input multiple-output (MIMO), minimum mean square error (MMSE), zero-forcing (ZF).
@@ I Introduction
In Part I of this paper, we have reviewed the growing interest in analog computing, driven by advancements in several technologies such as optical systems, metamaterials, and resistive memory arrays. Digital computing, despite its dominance, faces fundamental limitations in speed and power consumption, mainly due to the reliance on analog-to-digital converters (ADCs) and digital-to-analog converters (DACs). Analog computing addresses these challenges by enabling energy-efficient and highly parallelized computations at the speed of light. While Part I has focused on how analog computing can be leveraged for signal processing operations, Part II explores how these principles can be applied to achieve efficient and flexible beamforming in the analog domain, paving the way for future wireless communications.
@@
As we are witnessing a revival of interest in analog operations within the field of signal processing, a similar shift from digital to analog has emerged in wireless communications. This shift is driven by the growing number of antennas needed to support higher data rates and simultaneously serve a larger number of users. While massive multiple-input multiple-output (MIMO) systems typically deploy 64 or more antennas [2, 3], this number could reach several thousands in ultra-massive or gigantic MIMO systems, which have been recently proposed to meet the stringent requirements of 6G networks [4]. With such a large number of antennas, digital beamforming faces two critical challenges in terms of hardware and computational complexity. First, digital beamforming typically requires a dedicated radio-frequency (RF) chain per antenna element, each including ADCs/DACs and mixers, which are costly and power-hungry components. Second, as the number of antennas increases, the data volumes to be processed in real-time grow accordingly. For example, to precode the transmitted symbol vector at the transmitter (or combine the received symbol vector at the receiver) a matrix-vector product is required on a per-symbol basis. These two challenges result in high hardware costs and computational complexity when digital beamforming is applied in massive or gigantic MIMO systems. For these reasons, several alternatives to digital beamforming have been explored to enable beamforming operations fully or partially in the analog domain.
To address the high hardware cost associated with digital beamforming, alternative strategies such as analog beamforming and hybrid analog-digital beamforming have been proposed. Analog beamforming steers the beam pattern by solely requiring a single RF chain [5]. This is achieved by connecting the RF chain to the antennas via phase shifters or time delay elements, which impose a constant modulus constraint on the beamforming weights. Thus, while analog beamforming offers low hardware complexity, it is characterized by limited flexibility. To balance hardware cost and flexibility by bridging analog and digital beamforming, hybrid analog-digital beamforming techniques have been widely investigated [5, 6, 7, 8, 9]. In hybrid beamforming, the beamforming process is divided into two stages. First, the transmitted symbols are digitally processed in the baseband domain with a low-dimensional digital beamforming matrix, and fed to a reduced number of RF chains. Second, the RF chains are connected to the transmitting antennas through a network of phase shifters, operating analog beamforming in the RF domain. The reversed process is applied at the receiver.
@@
Another method to efficiently perform beamforming in the analog domain is given by reconfigurable intelligent surface (RIS) [10, 11]. This technology has recently emerged enabling the control of the wireless channel through surfaces made of multiple passive elements with reconfigurable scattering properties. Beamforming in the analog domain can be achieved through RIS by deploying a RIS in close proximity to an active transceiving device equipped with a reduced number of antennas. In this way, the effective radiation pattern of the active device can be optimized by reconfiguring the RIS elements [12, 13, 14]. Two extensions of this technology have been proposed to achieve additional flexibility over conventional RIS. First, beyond diagonal RIS (BD-RIS) include more general RIS architectures, characterized by a scattering matrix not restricted to be diagonal [15, 16, 17]. By deploying a BD-RIS close to an active device, its additional flexibility over conventional RIS in manipulating the radiation pattern can enable even higher performance [18]. Second, stacked intelligent metasurface (SIM) technology considers multiple stacked transmissive RISs deployed close to the active device [19, 20, 21]. The additional flexibility provided by the multiple layers has shown increased performance over just deploying a single conventional RIS.
In the reviewed beamforming strategies operating in the analog domain, the number of RF chains is commonly lower bounded by the number of symbols, or streams, transmitted in parallel. Nevertheless, other beamforming strategies have been developed to also perform symbol modulation in the analog domain, allowing the transmission of multiple symbols with a single RF chain. First, in electronically steerable parasitic array radiators (ESPARs), the symbols are modulated and the radiation pattern reconfigured by adjusting the mutual coupling between one active antenna element and multiple “parasitic” elements having tunable loads [22, 23, 24]. Second, in load modulated arrays (LMAs), a single RF source is connected to multiple antennas through tunable impedance components, which are reconfigured to achieve symbol modulation [25, 26]. Third, symbol modulation can be performed in the analog domain by employing a RIS deployed close to an active antenna transmitting a carrier signal [27, 28]. These three strategies further highlight the potential of leveraging analog-domain beamforming to reduce the hardware cost of transceiver devices.
In addition to alleviating the hardware cost, analog-domain strategies have been also explored to reduce the computational complexity of future wireless networks, by performing specific mathematical operations in the analog domain. To this end, in-memory computing has been proposed to accelerate ridge regression with applications to massive MIMO [29]. SIM has been used to compute the discrete Fourier transform (DFT) of the incident signal in the analog domain, which is particularly beneficial for direction-of-arrival (DoA) estimation [30]. Moreover, over-the-air computing emerged as a form of analog computing that leverages the superposition principle characterizing the electromagnetic (EM) propagation over the wireless channels. Over-the-air computing allows efficient computation of the so-called nomographic functions [31, 32, 33], with impactful applications on federated learning [34].
@@
Numerous technologies have been proposed to reduce the hardware cost of future wireless transceivers through analog beamforming [5]-[28]. However, operating in the analog domain typically decreases the flexibility compared to digital beamforming, with a consequent reduction in performance. Thus, whether it is possible to maintain maximum performance while simultaneously minimizing hardware cost and computational complexity through analog computing [29]-[34] remains an open question. In Part II of this paper, we address this gap by introducing a novel beamforming strategy enabled by microwave linear analog computers (MiLACs), denoted as MiLAC-aided beamforming. This beamforming strategy maintains the same flexibility as digital beamforming while operating entirely in the analog domain. Because of its fully analog nature, it minimizes the number of RF chains and relies only on low-resolution ADCs/DACs, reducing the hardware cost. By leveraging the computational capabilities of MiLACs, it also minimizes the computational complexity required for beamforming. More in detail, the contributions of Part II of this paper are as follows.
First, we demonstrate that a MiLAC can efficiently compute five mathematical operations of practical interest in MIMO communications, such as least squares (LS) and matched filtering (MF). To this end, we first show how these five operations can be seen as special cases of the linear minimum mean square error (LMMSE) estimator for linear observation processes, and then discuss how the microwave network of the MiLAC needs to be reconfigured to compute them. These operations can be computed with a MiLAC with significantly reduced computational complexity compared to conventional digital computing.
Second, we propose MiLAC-aided beamforming as a new beamforming strategy enabled by MiLAC scalable to gigantic MIMO systems. MiLAC-aided beamforming can perform precoding at the transmitter and combining at the receiver with arbitrary matrices, i.e., with the same flexibility as digital beamforming. While having the maximum flexibility, MiLAC-aided beamforming allows us to perform specific operations in the analog domain, such as zero-forcing (ZF) precoding at the transmitter and ZF combining at the receiver, with important gains in terms of computational complexity. In addition, it allows to precode or combine a signal with a fixed matrix, such as the DFT matrix, entirely in the analog domain, requiring no digital operations.
Third, we compare MiLAC-aided beamforming with existing beamforming strategies, namely digital, analog, hybrid, RIS-aided, and SIM-aided beamforming, in terms of flexibility of the beamforming matrix, hardware complexity, and computational complexity. As a result of this comparison, we discuss five benefits of MiLAC-aided beamforming: 1) it enables the maximum flexibility, 2) it requires the minimum number of RF chains, 3) it only needs low-resolution ADCs/DACs, 4) it does not require any computation on a per-symbol basis, and 5) it requires the minimum number of computations per coherence block for ZF.
@@
Fourth, we provide numerical results to assess the performance of MiLAC-aided beamforming in terms of sum rate and bit error rate (BER), confirming that it can achieve the same performance as digital beamforming. We also assess the benefits of MiLAC-aided beamforming over digital beamforming in terms of computational complexity, showing that it can perform ZF, MF, and the DFT with a gain in computational complexity of $1.5\times 10^{4}$, $2.0\times 10^{2}$, and $4.0\times 10^{7}$ times, respectively. Thus, in a gigantic MIMO system with thousands of antennas, MiLAC-aided beamforming achieves the same performance as digital ZF beamforming at just $1/15000$th of the computational cost. From a different perspective, a 4096-antenna MiLAC serving 4096 users via ZF beamforming requires the same computational complexity as a 256-antenna digital transmitter serving 256 users. This makes MiLAC-aided beamforming a game changer for gigantic MIMO, opening the door to radically new transceiver architectures that can massively scale the number of antennas and users.
Organization: In Section II, we briefly review the concept of MiLAC, as introduced in Part I of this paper. In Section III, we derive five special cases of the LMMSE estimator and show how they can be efficiently computed by a MiLAC. In Section IV, we propose MiLAC-aided beamforming as a novel beamforming strategy enabled by a MiLAC. In Section V, we compare MiLAC-aided beamforming with existing beamforming strategies. In Section VI, we provide numerical results to evaluate the performance and computational complexity of MiLAC-aided beamforming. Finally, Section VII concludes this paper.
Notation: Please, refer to the notation introduced in Section I of Part I of this paper.
@@ II Microwave Linear Analog Computer
In this section, we briefly review the concept of MiLAC, defined as an analog computer that linearly processes microwave signals, as introduced in Part I of this paper. In general, a MiLAC can be seen as a $P$-port reconfigurable microwave network, which can be characterized by its admittance matrix $\mathbf{Y}\in\mathbb{C}^{P\times P}$. We assume this microwave network to be made of $P^{2}$ tunable admittance components interconnecting each port to ground and to all the other ports, such that the microwave network can be reconfigured to assume any arbitrary admittance matrix $\mathbf{Y}$. The admittance component interconnecting port $k$ to ground is denoted as $Y_{k,k}\in\mathbb{C}$, for $k=1,\ldots,P$, while the admittance component interconnecting port $k$ to port $i$ is denoted as $Y_{i,k}\in\mathbb{C}$, for $\forall i\neq k$. The input signal $\mathbf{u}=[u_{1},\ldots,u_{N}]^{T}\in\mathbb{C}^{N\times 1}$ is applied by $N$ voltage generators with series admittance $Y_{0}$ to the first $N$ ports of the MiLAC, with $N\leq P$, as represented in Fig. 1. The output signal is read on all $P$ ports, where $P=N+M$ with $M$ denoting the number of ports with no input, and is denoted by $\mathbf{v}=[v_{1},\ldots,v_{P}]^{T}\in\mathbb{C}^{P\times 1}$, as shown in Fig. 1. For convenience of representation, we introduce the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ as $\mathbf{v}_{1}=[v_{1},\ldots,v_{N}]^{T}\in\mathbb{C}^{N\times 1}$ and $\mathbf{v}_{2}=[v_{N+1},\ldots,v_{p}]^{T}\in\mathbb{C}^{M\times 1}$, such that $\mathbf{v}=[\mathbf{v}_{1}^{T},\mathbf{v}_{2}^{T}]^{T}$. We consider a MiLAC with $M>0$ ports with no input in the following.
It has been shown in Part I of this paper that the output vectors $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$ can be expressed as a function of the admittance matrix $\mathbf{Y}$ and the input $\mathbf{u}$ by introducing a matrix $\mathbf{P}\in\mathbb{C}^{P\times P}$ as
@@
for $i,k=1,\ldots,P$, as discussed in Part I of this paper. Remarkably, a MiLAC can efficiently compute (3)-(4) and (5)-(6) in the analog domain, while they would require computationally expensive matrix-matrix products and matrix inversion operations if digitally computed.
III Analog Computing for the Special Cases of the LMMSE Estimator
In Part I of this paper, we have shown how a MiLAC can be used to compute the LMMSE estimator of a linear observation process with reduced computational complexity. In this section, we derive and discuss five special cases of the LMMSE estimator for linear observation processes, of particular interest in wireless communications. Then, we show how these five special cases can be efficiently computed by a MiLAC.
III-A LMMSE Estimator for Linear Observation Processes
Consider a linear observation process $\mathbf{y}=\mathbf{H}\mathbf{x}+\mathbf{n}$, where $\mathbf{y}\in\mathbb{C}^{Y\times 1}$ is the observation vector, $\mathbf{H}\in\mathbb{C}^{Y\times X}$ is a known constant matrix, $\mathbf{x}\in\mathbb{C}^{X\times 1}$ is the unknown random vector with mean $\bar{\mathbf{x}}=\mathbf{0}_{X\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{x}}$, and $\mathbf{n}\in\mathbb{C}^{Y\times 1}$ is the random noise vector with mean $\bar{\mathbf{n}}=\mathbf{0}_{Y\times 1}$ and covariance matrix $\mathbf{C}_{\mathbf{n}}$. We assume that $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ are known and invertible, and that the cross-covariance matrix of $\mathbf{x}$ and $\mathbf{n}$ is $\mathbf{C}_{\mathbf{x}\mathbf{n}}=\mathbf{0}_{X\times Y}$. As discussed in the part I of this paper, the LMMSE estimator of $\mathbf{x}$ is equivalently given by
where we have $\hat{\mathbf{x}}_{\textnormal{LMMSE},1}=\hat{\mathbf{x}}_{\textnormal{LMMSE},2}$ when $\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}+\mathbf{C}_{\mathbf{x}}^{-1}$ and $\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}+\mathbf{C}_{\mathbf{n}}$ are both invertible, which holds true in general.
III-B Special Cases of the LMMSE Estimator for Linear Observation Processes
The LMMSE estimator in Section III-A has been obtained with no assumptions on the covariance matrices $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ beyond their invertibility. With specific assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$, we can show that it includes five special cases of practical interest, as discussed in the following.
III-B1 Generalized Least Squares
In generalized least squares (GLS), also known as generalized linear regression, the variance on the a priori information on $\mathbf{x}$ is infinitely higher than on $\mathbf{n}$, i.e., $\|\mathbf{C}_{\mathbf{x}}\|_{F}\gg\|\mathbf{C}_{\mathbf{n}}\|_{F}$. Thus, $\mathbf{C}_{\mathbf{x}}^{-1}$ is negligible in (8) and, equivalently, $\mathbf{C}_{\mathbf{n}}$ is negligible in (9), yielding
$\hat{\mathbf{x}}_{\text{GLS},1}$ $=\left(\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{y},$ (10)
$\hat{\mathbf{x}}_{\text{GLS},2}$ $=\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\left(\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\right)^{-1}\mathbf{y},$ (11)
respectively. Note that, depending on the dimensions of $\mathbf{H}$, only one matrix among $\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{H}$ and $\mathbf{H}\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}$ may be invertible. Thus, only one among the estimates $\hat{\mathbf{x}}_{\text{GLS},1}$ and $\hat{\mathbf{x}}_{\text{GLS},2}$ may exist.
III-B2 Generalized Matched Filtering
Opposite to GLS, we can assume that the variance on the a priori information on $\mathbf{x}$ is infinitely smaller than on $\mathbf{n}$, i.e., $\|\mathbf{C}_{\mathbf{x}}\|_{F}\ll\|\mathbf{C}_{\mathbf{n}}\|_{F}$. In this case, that we denote as generalized matched filtering (GMF), both (8) and (9) boil down to
$\hat{\mathbf{x}}_{\text{GMF}}=\mathbf{C}_{\mathbf{x}}\mathbf{H}^{H}\mathbf{C}_{\mathbf{n}}^{-1}\mathbf{y}.$ (12)
III-B3 Regularized Least Squares
In regularized least squares (RLS), also known as ridge regression, the entries of $\mathbf{x}$ and $\mathbf{n}$ are uncorrelated and have equal variance, i.e., $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$ and $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, with $\lambda_{\mathbf{x}},\lambda_{\mathbf{n}}\in\mathbb{R}$ and $\lambda_{\mathbf{x}},\lambda_{\mathbf{n}}>0$. Thus, by introducing $\lambda\in\mathbb{R}$ such that $\lambda=\lambda_{\mathbf{n}}/\lambda_{\mathbf{x}}$, (8) and (9) simplify as
respectively. Interestingly, the expressions in (13) and (14) are widely used in MIMO wireless communications to realize the so-called regularized zero-forcing beamforming (R-ZFBF) and the minimum mean square error (MMSE) receiver, as effective linear transmission and reception techniques, respectively [35, Chapters 7, 12].
TABLE I: Special cases of the LMMSE estimator and their assumptions.
Assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$
LMMSE None
GLS $\|\mathbf{C}_{\mathbf{x}}\|_{F}\gg\|\mathbf{C}_{\mathbf{n}}\|_{F}$
GMF $\|\mathbf{C}_{\mathbf{x}}\|_{F}\ll\|\mathbf{C}_{\mathbf{n}}\|_{F}$
RLS $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$, $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$
OLS $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$, $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, $\lambda_{\mathbf{x}}\gg\lambda_{\mathbf{n}}$
OMF $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$, $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, $\lambda_{\mathbf{x}}\ll\lambda_{\mathbf{n}}$
III-B4 Ordinary Least Squares
In ordinary least squares (OLS), also simply known as LS, or linear regression, the entries of $\mathbf{x}$ and $\mathbf{n}$ are uncorrelated and have equal variance, i.e., $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$ and $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, and the variance on the a priori information on $\mathbf{x}$ is much higher than on $\mathbf{n}$, i.e., $\lambda_{\mathbf{x}}\gg\lambda_{\mathbf{n}}$. Thus, since $\lambda=\lambda_{\mathbf{n}}/\lambda_{\mathbf{x}}\ll 1$, (13) and (14) further simplify as
$\hat{\mathbf{x}}_{\text{OLS},1}$ $=\left(\mathbf{H}^{H}\mathbf{H}\right)^{-1}\mathbf{H}^{H}\mathbf{y},$ (15)
$\hat{\mathbf{x}}_{\text{OLS},2}$ $=\mathbf{H}^{H}\left(\mathbf{H}\mathbf{H}^{H}\right)^{-1}\mathbf{y},$ (16)
respectively. Interestingly, OLS is a special case of both GLS and RLS as it includes both their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$. As discussed for the GLS, since only one matrix among $\mathbf{H}^{H}\mathbf{H}$ and $\mathbf{H}\mathbf{H}^{H}$ may be invertible depending on the dimensions of $\mathbf{H}$, only one among $\hat{\mathbf{x}}_{\text{OLS},1}$ and $\hat{\mathbf{x}}_{\text{OLS},2}$ may exist, unless $\mathbf{H}$ is a square matrix. Remarkably, the expressions of the OLS are widely employed in MIMO wireless communications to implement zero-forcing beamforming (ZFBF) and the ZF receiver [35, Chapters 7, 12].
III-B5 Ordinary Matched Filtering
In ordinary matched filtering (OMF), also simply known as MF, the entries of $\mathbf{x}$ and $\mathbf{n}$ are uncorrelated and have equal variance, i.e., $\mathbf{C}_{\mathbf{x}}=\lambda_{\mathbf{x}}\mathbf{I}_{X}$ and $\mathbf{C}_{\mathbf{n}}=\lambda_{\mathbf{n}}\mathbf{I}_{Y}$, and the variance on the a priori information on $\mathbf{x}$ is much smaller than on $\mathbf{n}$, i.e., $\lambda_{\mathbf{x}}\ll\lambda_{\mathbf{n}}$. Thus, since $\lambda=\lambda_{\mathbf{n}}/\lambda_{\mathbf{x}}\gg 1$, both (13) and (14) boil down to
$\hat{\mathbf{x}}_{\text{OMF}}=\lambda^{-1}\mathbf{H}^{H}\mathbf{y}.$ (17)
Remarkably, OMF is a special case of both GMF and RLS as it jointly accounts for their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$. In addition, it is utilized in MIMO wireless communications to realize matched beamforming (MBF) and the MF receiver [35, Chapters 7, 12].
We summarize the five special cases of the LMMSE estimator for linear observation processes and their assumptions on $\mathbf{C}_{\mathbf{x}}$ and $\mathbf{C}_{\mathbf{n}}$ in Tab. I.
TABLE II: MiLAC output $\mathbf{v}_{2}$ for different $\mathbf{P}$ with $\mathbf{P}_{11}$ invertible.